If X is the number scored in a throw of a fair die, show that the Chebychev’s
inequality gives {P X −µ > }5.2 < 47.0 , where µis the mean of X, while the actual
probability is zero.
1
Expert's answer
2020-03-16T13:04:49-0400
"P\{|X-\mu|>2.5\}<0.47"


Chebyshev's inequality:


"P\{|X-\mu|\geq k\sigma\}\leq{1 \over k^2}, k\in \R"


Let "X" is our random variable.


"MX={1 \over 6}(1+2+3+4+5+6)={7 \over 2}"

"MX^2={1 \over 6}(1^2+2^2+3^2+4^2+5^2+6^2)={91 \over 6}"

"DX=MX^2-(MX)^2"

"DX={91 \over 6}-({7 \over 2})^2={35 \over 12}"

"\sigma_x=\sqrt{DX}=\sqrt{{35 \over 12}}"

"k\sigma_X=2.5=>k={2.5 \over \sqrt{{35 \over 12}}}=\sqrt{{15 \over 7}}\approx1.463850"

"{1 \over k^2}={7 \over 15}\approx0.4667<0.47"

Hence if "X" is the number scored in a throw of a fair die, then the Chebychev’s inequality gives 


"P\{|X-\mu|>2.5\}<0.47"